# Large deviations for the smallest eigenvalue of a deformed GOE with an outlier

### v1

@@ 2.4. A functional inequality for the large deviation upper bound
$\bullet$ If $\Phi(\Lambda,Y)\geq\omega_{\nu,t}(\ell_{\nu,t})$, then the smallest eigenvalue of $G_{N-1}+B_{N}^{\prime}$ converges towards $\ell_{\nu,t}$ and $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)$ goes to one as soon as $x<\ell_{\nu,t}$ and $\varepsilon$ is small enough.
@@
$\bullet$ If $x\leq H_{\nu,t}(\Phi(\Lambda,Y))$ and $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})$, then the smallest eigenvalue of $G_{N-1}+B_{N}^{\prime}$ goes to $H_{\nu,t}(\Phi(\Lambda,Y))$ when $N$ goes to infinity and is therefore greater than $x-2\varepsilon$. Hence, we see that $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)$ goes to one as $N$ goes to infinity.
@@
$\bullet$ If $x>H_{\nu,t}(\Phi(\Lambda,Y))$ and $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})$, then $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)$ goes to zero exponentially fast by large deviations. We can replace $B_{N}^{\prime}$ by another matrix $B_{N}^{\prime\prime}$ with the same eigenvalues except that $\Lambda_{1}$ is replaced by $\Phi(\Lambda,Y)$ up to an additional error of order $\varepsilon$. Observe that $\lambda_{1}(G_{N-1}+B_{N}^{\prime\prime})\leq\ell_{\nu,t}$ with overwhelming probability by (2.20). We can cut the compact set $[x-3\varepsilon,\ell_{\nu,t}]$ into a finite number $K$ of intervals of size $\gamma$ and replace $\Lambda_{1}$ by $\Phi(\Lambda,Y)$ up to a small error (an error of size $2\varepsilon$) to deduce that
$\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)\leq\sum_{i=1}^{K}P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-3\varepsilon+i\gamma,\gamma)\,.$
@@
$P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-3\varepsilon+i\gamma,\gamma)\leq 2N^{2}P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-3\varepsilon,\gamma)+2N^{2}e^{-cN^{2}}$
@@
$\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)\leq 2KN^{2}P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-3\varepsilon,\gamma)+2N^{2}Ke^{-cN^{2}}\,.$
@@
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon)\leq-F^{+}_{\nu,t}(\Phi(\Lambda,Y),x-3\varepsilon)\,.$
To conclude we need to first take the supremum over $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$, so $\mathsf{M}^{\prime}\in\mathcal{B}^{N}_{\delta(\alpha)}$ before the large $N$ limit. We have to be careful because $\nu$ depends implicitly on $\mathsf{M}/N$ and $Y_{N}$ is not equal to $Y$ but is close to it. But the condition $\{\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})\}\cap\{x\geq H_{\nu,t}(\Phi(\Lambda,Y))\}=\{\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y)\}$ is continuous in $\nu$ (as $\omega_{\nu,t}(x)$ is) and $Y$. Moreover, $F^{+}_{\nu,t}(\Phi(\Lambda,Y),x-3\varepsilon)$ is lower semicontinuous in $Y$. The argument above is therefore valid uniformly on $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$ as long as we are in the interior of the set $\{\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y)\}$ for $\delta$ small enough. We therefore conclude that for $\kappa$ small enough,
$\lim_{\delta\downarrow 0}\limsup_{N\to\infty}\sup_{v_{1}\in E_{Y_{k},\varepsilon^{\prime}}\atop M\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log\Bigr{(}\frac{1}{Z_{N-1}^{t\frac{N-1}{N}}}\int_{x-\varepsilon\leq\lambda_{2}\leq\cdots\leq\lambda_{N}}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B_{N})p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}\Bigr{)}$
$\qquad\leq-F_{\nu,t}^{+}(\Phi(\Lambda,Y_{k}),x-3\varepsilon)1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y_{k})}+o(\kappa,\varepsilon).$
@@
We deduce from (2.23) that
$\lim_{\delta\downarrow 0}\limsup_{N\to\infty}\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\\ \leq\sup_{k}\Bigr{\{}-L_{\nu,t}^{\Lambda}(\lambda,Y_{k})-I_{\nu}(Y_{k})-F_{\nu,t}^{+}(\Phi(\Lambda,Y_{k}),x-3\varepsilon)1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y_{k})}\Bigr{\}}\\ +\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)-\frac{\lambda^{2}}{4t}-C_{t}+o(\varepsilon^{\prime},\kappa,\varepsilon).\$
@@
$\lim_{\varepsilon\downarrow 0}\limsup_{N\to\infty}\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\\ \leq\sup_{k}\Bigr{\{}-K_{\nu,t}(\Lambda,x,Y_{k})-F_{\nu,t}^{+}(\Phi(\Lambda,Y_{k}),x^{-})1_{\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y_{k})}\Bigr{\}}+o(\varepsilon^{\prime},\kappa)\\ \leq\sup_{Y\in\Delta}\Bigr{\{}-K_{\nu,t}(\Lambda,x,Y)-F_{\nu,t}^{+}(\Phi(\Lambda,Y),x^{-})1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y)}\Bigr{\}}+o(\varepsilon^{\prime},\kappa).$
We can finally let $\varepsilon^{\prime}$ going to zero. Now, we take the supremum over $\Lambda\in[\Lambda^{\prime}-\varepsilon,\Lambda^{\prime}+\varepsilon]$ on both sides. In the right-hand side it can be interchanged with the supremum over $Y$. The map $J_{\nu,t}^{\Lambda}(\lambda,Y)=L_{\nu,t}^{\Lambda}(\lambda,Y)+I_{\nu}(Y)$ is uniformly continuous in $\Lambda$. Moreover
@@ 2.5. A functional inequality for the large deviation lower bound
$F_{\nu}^{-}(\Lambda,\lambda):=-\liminf_{{\Lambda^{\prime}\to\Lambda}}\liminf_{\varepsilon\to 0}\lim_{\delta\downarrow 0}\liminf_{N\to\infty}\inf_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda^{\prime},\mathsf{M}}(\lambda,\varepsilon).$ (2.24)
@@
Let $\Lambda\leq\ell_{\nu}$. For every $\lambda<\ell_{\nu,t}$, we have
@@
Again $B_{N}$ has eigenvalues $\eta_{i}$ with multiplicity $M_{i}$ and $\Lambda^{\prime}$ with multiplicity $1$. Let $Y\in\Delta^{p+1}$ and $\kappa,\varepsilon^{\prime}>0$. Letting $E_{Y,\kappa,\varepsilon^{\prime}}$ be as in (2.22), one can bound $P_{N,\Lambda^{\prime},\mathsf{M}}(x,\varepsilon)$ from below by
$P_{N,\Lambda^{\prime},\mathsf{M}}(x,\varepsilon)\geq\frac{1}{Z_{N}^{t}}\int_{E_{Y,\kappa,\varepsilon^{\prime}}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(v_{2},\ldots,v_{N})\\ \int_{\mathcal{B}(\delta,\varepsilon,x)}e^{-NL_{\nu,t}^{\Lambda^{\prime}}(\lambda_{1},\langle v_{a},w_{0}\rangle^{2},...)-\frac{N}{4t}\lambda_{1}^{2}}\prod_{1\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B)p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{1}\cdots\mathrm{d}\lambda_{N},$ (2.26)
@@
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\geq\frac{2\varepsilon Z^{t(1-N^{-1})}_{N-1}}{Z^{t}_{N}}e^{N(\int\log|x-\lambda|\mathrm{d}(\sigma_{t}\boxplus\nu)(\lambda)-L_{\nu,t}^{\Lambda}(x,Y)+o(\delta,\epsilon)-\frac{1}{4t}x^{2}+o(\varepsilon,\delta))}\\ \times\int_{E_{Y,\kappa,\varepsilon^{\prime}}}\mathrm{d}v_{1}\mathbb{P}(\lambda(G_{N-1}+p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp})\in C(\delta,\varepsilon,x)).$
Moreover, recall that $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$, as a $(N-1)\times(N-1)$ matrix in $v_{1}^{\perp}$, has a unique outlier in $[\Lambda,\ell_{\nu}]$, which is close to $\Phi(\Lambda^{\prime},Y)$, and eigenvalues $\eta_{i}$ with multiplicities in $\{N_{i},N_{i}-1\}$. Let $B_{N}^{\prime}$ be the matrix with outlier $\Phi(\Lambda^{\prime},Y)$ and eigenvalues $\eta_{i}$ with multiplicity $M_{i}^{\prime}=M_{i}$ except for $\eta_{p}$ with multiplicity $M_{p}^{\prime}=M_{p}-1$ (with the same eigenspaces as $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$). Then $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}\geq B_{N}^{\prime}$ so that
$\mathbb{P}(\lambda_{1}(G_{N-1}+p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp})\in C(\delta,\varepsilon,x))$$\geq$ $\mathbb{P}(\lambda_{1}(G_{N-1}+p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp})\geq x+2\varepsilon)$
@@
$\geq$ $\mathbb{P}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x+2\varepsilon)-e^{-cN^{2}},$
@@
where we took $\delta$ to zero. We can then let $\varepsilon$ going to zero and take the supremum over all possible $Y$. We can finally take the infimum over $\Lambda\in[\Lambda^{\prime}-\varepsilon,\Lambda^{\prime}+\varepsilon]$ and use the upper semi-continuity of $F_{\nu,t}^{-}$. This completes the proof of (2.25) as soon as $x+3\varepsilon<\ell_{\nu,t}$, which is true for $\varepsilon$ small enough as long as $x<\ell_{\nu,t}$. ∎
2.6. Solving the functional equation in the region $\omega_{\nu,t}(\lambda)\leq\Lambda$

### v2

@@ 2.4. A functional inequality for the large deviation upper bound
$\bullet$ If $\Phi(\Lambda,Y)\geq\omega_{\nu,t}(\ell_{\nu,t})$, then the smallest eigenvalue of $G_{N-1}+B_{N-1}^{\prime}$ converges towards $\ell_{\nu,t}$ and $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N-1}^{\prime\prime})\geq x-3\varepsilon)$ goes to one as soon as $x<\ell_{\nu,t}$ and $\varepsilon$ is small enough.
@@
$\bullet$ If $x\leq H_{\nu,t}(\Phi(\Lambda,Y))$ and $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})$, then the smallest eigenvalue of $G_{N-1}+B_{N-1}^{\prime\prime}$ goes to $H_{\nu,t}(\Phi(\Lambda,Y))$ when $N$ goes to infinity and is therefore greater than $x-3\varepsilon$. Hence, we see that $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N-1}^{\prime\prime})\geq x-2\varepsilon)$ goes to one as $N$ goes to infinity.
@@
$\bullet$ If $x>H_{\nu,t}(\Phi(\Lambda,Y))$ and $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})$, then $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N-1}^{\prime\prime})\geq x-3\varepsilon)$ goes to zero exponentially fast by large deviations. We can replace $B_{N}^{\prime}$ by another matrix $B_{N}^{\prime\prime}$ with the same eigenvalues except that $\Lambda_{1}$ is replaced by $\Phi(\Lambda,Y)$ up to an additional error of order $\varepsilon$. Observe that $\lambda_{1}(G_{N-1}+B_{N}^{\prime\prime})\leq\ell_{\nu,t}$ with overwhelming probability by (2.23). We can cut the compact set $[x-4\varepsilon,\ell_{\nu,t}]$ into a finite number $K$ of intervals of size $\gamma$ and replace $\Lambda_{1}$ by $\Phi(\Lambda,Y)$ up to a small error (an error of size $\varepsilon$) if $\kappa$ is small enough to deduce that
$\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime\prime})\geq x-3\varepsilon)\leq\sum_{i=1}^{K}P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-4\varepsilon+i\gamma,\gamma)\,.$
@@
$P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-4\varepsilon+i\gamma,\gamma)\leq 2N^{2}P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-4\varepsilon,\gamma)+2N^{2}e^{-cN^{2}}$
@@
$\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime\prime})\geq x-3\varepsilon)\leq 2KN^{2}P_{N-1,\Phi(\Lambda,Y),\mathsf{M}^{\prime}}(x-4\varepsilon,\gamma)+2N^{2}Ke^{-cN^{2}}\,.$
@@
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime\prime})\geq x-3\varepsilon)\leq-F^{+}_{\nu,t}(\Phi(\Lambda,Y),x-4\varepsilon)\,.$
To conclude we need to first take the supremum over $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$, so $\mathsf{M}^{\prime}\in\mathcal{B}^{N}_{\delta(\alpha)}$ before the large $N$ limit. We have to be careful because $\nu$ depends implicitly on $\mathsf{M}/N$ and $Y_{N}$ is not equal to $Y$ but is close to it. But the condition $\{\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\ell_{\nu,t})\}\cap\{x\geq H_{\nu,t}(\Phi(\Lambda,Y))\}=\{\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y)\}$ is continuous in $\nu$ (as $\omega_{\nu,t}(x)$ is) and $Y$. Moreover, $F^{+}_{\nu,t}(\Phi(\Lambda,Y),x-4\varepsilon)$ is lower semicontinuous in $Y$. The argument above is therefore valid uniformly on $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$ as long as we are in the interior of the set $\{\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y)\}$ for $\delta$ small enough. We therefore conclude that for $\kappa$ small enough,
$\lim_{\delta\downarrow 0}\limsup_{N\to\infty}\sup_{v_{1}\in E_{Y_{k}}\atop M\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log\mathbb{K}(x,\varepsilon,v_{1})\leq-F_{\nu,t}^{+}(\Phi(\Lambda,Y_{k}),x-4\varepsilon)1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y_{k})}+o(\kappa).$ (2.31)
@@
We therefore deduce from (2.27), (2.30), (2.31) that
$\lim_{\delta\downarrow 0}\limsup_{N\to\infty}\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\\ \leq\max\Big{\{}\sup_{k}\Bigr{\{}-L_{\nu,t}^{\Lambda}(x,Y_{k})-I_{\nu}(Y_{k})-F_{\nu,t}^{+}(\Phi(\Lambda,Y_{k}),x-4\varepsilon)1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y_{k})}\Bigr{\}}\\ +\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(\lambda)-\frac{x^{2}}{4t}-C_{t}+o(\varepsilon^{\prime},\kappa,\varepsilon),\frac{1}{2}\alpha_{1}\log\zeta/\alpha_{1}+C^{\prime}\Big{\}}.\$
@@
$\lim_{\varepsilon\downarrow 0}{\lim_{\delta\downarrow 0}}\limsup_{N\to\infty}\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\\ \leq\max\{\sup_{k}\Bigr{\{}-K_{\nu,t}(\Lambda,x,Y_{k})-F_{\nu,t}^{+}(\Phi(\Lambda,Y_{k}),x^{-})1_{\omega_{\nu,t}(x)\geq\Phi(\Lambda,Y_{k})}\Bigr{\}}+o(\varepsilon^{\prime},\kappa),\frac{1}{2}\alpha_{1}\log\frac{\zeta}{\alpha_{1}}+C^{\prime}\}\\ \leq\max\{\sup_{Y\in\Delta}\Bigr{\{}-K_{\nu,t}(\Lambda,x,Y)-F_{\nu,t}^{+}(\Phi(\Lambda,Y),x^{-})1_{\omega_{\nu,t}(x)>\Phi(\Lambda,Y)}\Bigr{\}}+o(\varepsilon^{\prime},\kappa),\frac{1}{2}\alpha_{1}\log\frac{\zeta}{\alpha_{1}}+C^{\prime}\}.$
We can finally let $\varepsilon^{\prime}$, $\kappa$ going to zero and finally $\zeta$ to zero. Now, we take the supremum over $\Lambda\in[\Lambda^{\prime}-\varepsilon,\Lambda^{\prime}+\varepsilon]$ on both sides. In the right-hand side it can be interchanged with the supremum over $Y$. The map $J_{\nu,t}^{\Lambda}(\lambda,Y)=L_{\nu,t}^{\Lambda}(\lambda,Y)+I_{\nu}(Y)$ is uniformly continuous in $\Lambda$. Moreover
@@ 2.5. A functional inequality for the large deviation lower bound
$F_{\nu}^{-}(\Lambda,\lambda):=-\liminf_{{\Lambda^{\prime}\to\Lambda}}\lim_{\varepsilon\downarrow 0}\lim_{\delta\downarrow 0}\liminf_{N\to\infty}\inf_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda^{\prime},\mathsf{M}}(\lambda,\varepsilon).$ (2.32)
@@
Let $\Lambda\leq\ell_{\nu}$. Let $\Phi(\Lambda,\cdot)$ be as in (2.14) and $\ell_{\nu,t}^{\Lambda}$ be as in (1.14). For every $\lambda<\ell_{\nu,t}$, we have
@@
Again $B_{N}$ has eigenvalues $\eta_{i}$ with multiplicity $M_{i}$ and $\Lambda^{\prime}$ with multiplicity $1$. Let $Y\in\Delta^{p+1}$ and $\kappa,\varepsilon^{\prime}>0$. We may and shall assume that $y_{1}>0$ since we saw in the upper bound that small $y_{1}$ have exponentially small contributions. We also can take $\kappa$ small enough so that $y_{1}$ is bounded below by a positive constant uniformly on $E_{Y,\kappa,\varepsilon^{\prime}}$ defined in (2.26). One can bound $P_{N,\Lambda^{\prime},\mathsf{M}}(x,\varepsilon)$ from below by
$P_{N,\Lambda^{\prime},\mathsf{M}}(x,\varepsilon)\geq\frac{1}{Z_{N}^{t}}\int_{E_{Y,\kappa,\varepsilon^{\prime}}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(v_{2},\ldots,v_{N})\\ \int_{\mathcal{B}(\delta,\varepsilon,x)}e^{-NL_{\nu,t}^{\Lambda^{\prime}}(\lambda_{1},\langle v_{a},w_{0}\rangle^{2},...)-\frac{N}{4t}\lambda_{1}^{2}}\prod_{1\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B)p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda^{N},$ (2.34)
@@
$P_{N,\Lambda^{\prime},\mathsf{M}}(x,\varepsilon)\geq\frac{2\varepsilon Z^{t(1-N^{-1})}_{N-1}}{Z^{t}_{N}}e^{N(\int\log|x-\lambda|\mathrm{d}(\sigma_{t}\boxplus\nu)(\lambda)-L_{\nu,t}^{\Lambda^{\prime}}(x,Y)+o(\delta,\epsilon)-\frac{1}{4t}x^{2}+o(\varepsilon,\delta))}\\ \times\int_{E_{Y,\kappa,\varepsilon^{\prime}}}\mathrm{d}v_{1}\mathbb{P}(\lambda(G_{N-1}+B_{N-1}^{\prime})\in C(\delta,\varepsilon,x)).$
Moreover, recall that $B_{N-1}^{\prime}$, as a $(N-1)\times(N-1)$ matrix in $v_{1}^{\perp}$, has a unique outlier in $[\Lambda,\ell_{\nu}]$, which is close to $\Phi(\Lambda^{\prime},Y)$, and eigenvalues $\eta_{i}$ with multiplicities in $\{N_{i},N_{i}-1\}$. Let $B_{N-1}^{\prime\prime}$ be the matrix with outlier $\Phi(\Lambda^{\prime},Y)$ and eigenvalues $\eta_{i}$ with multiplicity $M_{i}^{\prime}=M_{i}$ except for $\eta_{p}$ with multiplicity $M_{p}^{\prime}=M_{p}-1$ (with the same eigenspaces as $B_{N-1}^{\prime}$). Then $B_{N-1}^{\prime}\geq B_{N-1}^{\prime\prime}$ so that
$\mathbb{P}(\lambda(G_{N-1}+B_{N-1}^{\prime})\in C(\delta,\varepsilon,x))$$\geq$ $\mathbb{P}(\lambda_{1}(G_{N-1}+B_{N-1}^{\prime})\geq x+2\varepsilon)$
@@
$\geq$ $\mathbb{P}(\lambda_{1}(G_{N-1}+B_{N-1}^{\prime\prime})\geq x+2\varepsilon)-e^{-cN^{2}},$
@@
where we took $\delta$ to zero. We can then let $\varepsilon$ going to zero and take the supremum over all possible $Y$. We can finally take the infimum over $\Lambda^{\prime}\in[\Lambda-\varepsilon,\Lambda+\varepsilon]$ and use the upper semi-continuity of $F_{\nu,t}^{-}$. This completes the proof of (2.33) as soon as $x+3\varepsilon<\ell_{\nu,t}$, which is true for $\varepsilon$ small enough as long as $x<\ell_{\nu,t}$. ∎
2.6. Solving the functional equation in the complement of the region $\Lambda<\omega_{\nu,t}(\lambda)$